OpenStudy (anonymous):

f(x) varies inversely with x and f(x) = -10 when x=5. What is the inverse variation equation?

OpenStudy (trojanpoem):

f(x) = k/x -10 = k/5 k = - 50 equation = f(x) = -50/x Try to do them your self